If someone can please help with this question, because I don't understand how to solve it!
Mr. Moore is 6 years older than 4 times his daughters' age. In 7 years, he will be triple her age. How old will Mr. Moore be in 9 years ?
What is his daughters age now ?

Answer by mananth(16946)

If someone can please help with this question, because I don't understand how to solve it!
Mr. Moore is 6 years older than 4 times his daughters' age. In 7 years, he will be triple her age. How old will Mr. Moore be in 9 years ?
What is his daughters age now ?
let daughter's age =x
daughter-------------------- Moore
x----------------------------4x+6
after seven years ( add seven years to their present ages)
Moore is 3 times her age
(x+7)*3--------------------- (4x+6)+7
(x+7)*3 = (4x+6)+7
3x+21=4x+13
4x-3x=21-13
x=8 daughter's age = 8
Moore's age after 7 years will be
(4x+6)+7
38+7= 45is moore's age after 7 years
add2 to it to get Moore's age after 9 years = 47